### v1

# Text-to-SQL Task-oriented Dialogue Ontology Construction

Abstract
Large language models (LLMs) are widely used as general-purpose knowledge sources, but they rely on parametric knowledge, limiting explainability and trustworthiness. In task-oriented dialogue (TOD) systems, this separation is explicit, using an external database structured by an explicit ontology to ensure explainability and controllability. However, building such ontologies requires manual labels or supervised training.
We introduce TeQoDO: a Text-to-SQL task-oriented Dialogue Ontology construction method. Here, an LLM autonomously builds a TOD ontology from scratch without supervision using its inherent SQL programming capabilities combined with dialogue theory provided in the prompt.
We show that TeQoDO outperforms transfer learning approaches, and its constructed ontology is competitive on a downstream dialogue state tracking task. Ablation studies demonstrate the key role of dialogue theory. TeQoDO also scales to allow construction of much larger ontologies, which we investigate on a Wikipedia and ArXiv dataset. We view this as a step towards broader application of ontologies to increase LLM explainability. [^1] Code will be released upon publication.
1 Introduction
Large language models (LLMs) have become ubiquitous knowledge processing systems, with the ability to reach or surpass human performance on a wide variety of natural language processing tasks. These models are pre-trained on massive corpora and aligned via human feedback using reinforcement learning Brown et al. (2020); Ouyang et al. (2022). Despite these remarkable abilities, there are some inherent problems associated with these systems. Their knowledge is stored in vast numbers of parameters, which makes it very difficult to understand their behaviour, even via probing Cífka and Liutkus (2023). It is therefore extremely challenging to verify their inherent knowledge Zhong et al. (2024). Moreover, LLMs often produce confident but non-factual outputs that just appear plausible on a superficial level Sahoo et al. (2024); Feng et al. (2024). Finally, many LLMs and their training data are closed-source.
Ontologies provide a human-readable means of reconstructing knowledge-based reasoning in language models Gruber (1995); Lo et al. (2024). Manually building ontologies for all relevant domains is infeasible Milward and Beveridge (2003). To ensure broad applicability, ontologies should be constructed automatically, reliably, and consistently. Automatic ontology construction presents a promising solution to these challenges.
In this work, we focus on constructing task-oriented dialogue (TOD) ontologies from raw dialogue data. The goal is to extract relevant information and organise it into a meaningful hierarchy for handling user queries. TOD ontologies consist of domains, slots, and values, with system actions and user intents defined over this domain-slot-value structure. Figures˜1 and 2 illustrate ontology construction in this setting. Such ontologies are essential for modern TOD systems Young et al. (2013); Hudeček and Dusek (2023).
Most TOD ontology construction approaches follow two steps Hudeček et al. (2021); Vukovic et al. (2022): (1) term extraction from dialogue data and (2) relation extraction between the terms from the previous step. Existing approaches tackle these steps separately and rely on annotated training data Vukovic et al. (2024); Finch et al. (2024). This has the downside of additional and potentially error-prone training, and the danger of information loss between the two processing steps.
We propose TeQoDO, a Text-to-SQL task-oriented Dialogue Ontology construction approach. Figure˜3 provides an overview of the method. TeQoDO uses LLMs’ code understanding and generation capabilities Chen et al. (2023) to build ontologies from scratch via SQL. It incrementally constructs the database by iterating over dialogues, retrieving relevant existing content, and updating the DB with new information. The model uses state tracking from dialogue theory to distinguish new content from existing data in the DB. The update prompt includes a notion of success to help align schema changes with the user’s goals. Query results are augmented with semantically similar concepts or example values from existing columns to improve coverage and consistency.
Text-to-SQL provides a structured format familiar to LLMs through code encountered during pre-training Deng et al. (2022); Zhao et al. (2024). It reduces the need for textual prompts, as SQL tables, columns, and values align with TOD ontology domains, slots, and values. To our knowledge, we are the first to use text-to-SQL to build a TOD ontology from scratch. We also adapt similarity-based evaluation metrics for ontology learning Lo et al. (2024) to reflect the hierarchical nature of TOD ontologies. This enables evaluation of various concept types, such as domains and slots.
We run experiments on two widely used TOD datasets: MultiWOZ Eric et al. (2020) and Schema-Guided Dialogue (SGD) Rastogi et al. (2020). TeQoDO outperforms recent TOD ontology construction methods Finch et al. (2024); Vukovic et al. (2024). To test generalisation, we apply TeQoDO to general ontology datasets from Lo et al. (2024) and show it scales to large ontologies. In summary, our contributions are:
  • We propose a text-to-SQL framework for building TOD ontologies from scratch, allowing LLMs to update a DB with each newly observed dialogue without supervision or rule-based aggregation.
  • The constructed ontology enables dialogue state tracking with performance comparable to using the ground truth ontology.
  • TeQoDO generalises to large-scale ontology datasets like Wikipedia and ArXiv, performing competitively.
2 Related Work
2.1 Ontology Construction
Figure 1: An example of ontology construction shows the extraction of the domain “Hotels”, with the value “cheap” assigned to the slot “Price”. System actions and user intents are defined based on the domains and slots specified in the ontology.
Prior to deep learning, ontology construction relied on frequency-based rules and linguistic features. With language models, clustering and supervised methods enabled feature-based learning, as outlined below.
Rule-based
ontology construction (Frantzi and Ananiadou, 1999; Nakagawa and Mori, 2002) relies on frequency-based methods to extract key subsequences from text. Wermter and Hahn (2006) use linguistic features for more advanced term extraction. While interpretable, these methods are hard to adapt across domains and often yield low precision.
Clustering-based
approaches, such as Yu et al. (2022), apply unsupervised parsing and hierarchical clustering to group extracted terms into domains and slots. Finch et al. (2024) train a generative slot induction model to extract domain-slot-value triples, which are then clustered into a slot hierarchy. These methods usually produce many slots, whose interpretation depends entirely on how well they match the ground truth. This is because evaluation compares slots solely based on their values. They also rely heavily on data quality, embedding representations, and sensitive hyperparameters, such as the minimum cluster size.
Supervised
Lo et al. (2024) fine-tune LLMs to predict ontology sub-graphs at the document level and merge them using frequency-based rules to construct ontologies for Wikipedia and ArXiv datasets. They also propose evaluation functions covering both verbatim content and higher-level structural semantics compared to the ground truth, which we adapt to TOD. Vukovic et al. (2024) train a model to predict TOD ontology subgraphs at the dialogue level and aggregate them by merging all predictions. They improve generalisability by updating the decoding with constraints and confidence-based adjustments. However, their method focuses only on relation extraction, using ground truth terms as input. These methods require costly annotated training data from Budzianowski et al. (2018).
In contrast, our approach requires no training and conducts both term and relation extraction. We use Vukovic et al. (2024) and Finch et al. (2024) as competitive baselines.
2.2 Text-to-SQL
Li et al. (2024) present a large benchmark for text-to-SQL queries on big databases and find LLMs underperform compared to humans. Zhou et al. (2024) apply LLMs to database tasks such as query rewriting and index tuning using automatic prompt generation, DB-specific pre-training, model design, and fine-tuning. Yang et al. (2024) enhance text-to-SQL parsing for smaller models by synthesising training data with larger models. Allemang and Sequeda (2024) improve text-to-SPARQL Polleres (2014) by using a rule-based query checker leveraging the underlying ontology, with an LLM repairing queries based on this feedback. In contrast to our work, these methods work with existing ontologies and do not build ontologies from scratch using text-to-SQL.
3 Text-to-SQL Ontology Construction
Figure 2: Example TOD ontology.
Figure 3: TeQoDO Overview with example DB queries and results.
3.1 Problem Formulation
Given a task-oriented dialogue dataset $D=\{d_{1},\ldots,d_{k}\}$, each dialogue $d_{i}$ is comprised of alternating user and system turns $\{u_{1},s_{1},\ldots,u_{m_{i}},s_{m_{i}}\}$ containing unstructured information about user-queried entities. The goal is to induce an ontology $O_{D}$ for the dataset. The ontology is a directed graph $O_{D}=(V,E)$ with five node types $V=V_{\text{domain}}\cup V_{\text{user\_intent}}\cup V_{\text{system\_action}}\cup V_{\text{slot}}\cup V_{\text{value}}$, forming a three-level hierarchy. The edge set $E$ is a subset of edges between specific node types, $E\subseteq(V_{\mathrm{domain}}\times V_{\mathrm{slot}})\cup(V_{\mathrm{slot}}\times V_{\mathrm{value}})\cup(V_{\mathrm{user\_intent}}\times V_{\mathrm{user\_intent}})\cup(V_{\mathrm{system\_action}}\times V_{\mathrm{system\_action}}).$
See Figure˜2 for an example TOD ontology, with domains in blue, slots in brown, values in green, and intents and actions in red. Domains represent broad topics, while slots specify particular types of information. Values provide concrete content for the slots. User intents and system actions define how domain-slot pairs can be used, based on the domain-slot-value hierarchy.
3.2 Method
Our main prompting approach combines task-oriented dialogue modelling with SQL, as outlined below. We then detail the steps of our iterative ontology construction pipeline. We utilise an LLM that creates SQL queries, which are then executed using Python.
SQL-Background
SQL is a query language for relational databases Chamberlin and Boyce (1974). We use SQLite, [^2] https://www.sqlite.org a serverless relational DB management system suitable for small-scale databases and schema-based knowledge representation. SQLite stores data in tables with columns and values, e.g., a “restaurant” table with a “price” column and “expensive” as a value. This structure aligns with TOD ontology formats, making it well-suited for ontology construction. For intents and actions, there are separate tables that store names of intent and action entries.
TeQoDO primarily uses data retrieval, definition, and manipulation queries. To access current DB content, a SELECT query retrieves table names. PRAGMA queries fetch column names and data types for each table. SELECT can also query specific values from tables and columns. Tables are created with CREATE TABLE, specifying column names and types. ALTER TABLE adds columns, INSERT INTO adds entities, and UPDATE modifies column values.
Task-oriented dialogue modelling
We incorporate two concepts from modular task-oriented dialogue models (Young et al., 2013) into our prompt to improve the quality of generated update queries. First, dialogue state tracking (DST) is used as a separate step to distinguish existing database content from new input, based on the current schema. Second, we prompt the model to consider dialogue success, guiding it to make updates that support achieving the user goal. For a more formal description, see Table˜7 in the appendix.
TeQoDO Prompting Steps
See Figure˜3 for an overview of TeQoDO with example queries and results. More detailed pipeline steps are given in Algorithm˜1 and are described in the following paragraphs.
Step 1: Query existing DB information
To ensure structural and naming consistency in TeQoDO, the model is first prompted to query the current DB tables. Initially, as the ontology is built from scratch, the DB is empty, and thus no results are returned. The resulting table list $T=\{t_{1},\ldots,t_{n}\}$ is included with the dialogue in step 2. The model is prompted to query column info for tables relevant to the dialogue. We then append the column query results to the prompt so the model generates SELECT queries consistent with the DB state. In step 3, it generates SELECT queries to retrieve specific entity values from the dialogue.
Step 2: Dialogue State Tracking with DB information
In this step, the model restructures DB query results into dialogue state tracking style labels. It predicts the current dialogue’s value for each table-column pair and compares it with the DB’s stored information. This improves the distinction between existing DB information and new information from the dialogue.
The model is prompted to summarise dialogue states using the DB’s table and column names, as well as intents and actions. This summary includes information from both the user and the system. By predicting the state, the model condenses the dialogue’s information into a concise, model-determined format not fully detailed in the prompt. Notably, the model tracks the entire dialogue at once, not turn-by-turn. It must track all mentioned information, not only what the user queries.
Step 3: DB Update
In the final step, the model generates database update queries based on the current DB state and dialogue information. It is prompted to create update queries considering DB query and DST results. Update queries include CREATE TABLE, ALTER TABLE to add columns, and UPDATE to add values. The model must follow the existing DB structure and generate consistent updates, preferring existing tables over creating new ones. Additionally, the model is prompted to update the DB to support fulfilling the user’s goal by incorporating dialogue success into the prompt.
4 Experiments
4.1 Datasets
MultiWOZ
The first dataset we employ is MultiWOZ 2.1 (Eric et al., 2020), a large-scale, multi-domain dialogue dataset containing human-human conversations annotated with domain, slot, and value labels. It covers information such as hotel bookings, restaurant reservations, taxi services, and attractions. We utilise the test set with 1,000 dialogues and 6 domains.
Algorithm 1 TeQoDO
1Input: Dialogue dataset $D$, Existing Database DB, Prompt $p_{0}$
2for $d_{i}$ in $D$ do
3  Query current set of tables $T=\{t_{1},...,t_{n}\}\in\textup{DB}$:
4  Query table columns $c_{T,d_{i}}$ using prompt $p=p_{0}+d_{i}+T$
5  Generate SELECT queries $v_{T,d_{i}}$ using $p=p+c_{T,d_{i}}$
6  Track information $ST_{d_{i}}$ using $p=p+v_{T,d_{i}}$
7  Generate update queries $U_{d_{i}}$ for success using $p=p+ST_{d_{i}}$
8end for
SGD
Second, we use the schema-guided dialogue (SGD) dataset (Rastogi et al., 2020). It is a diverse, multi-domain dataset for TOD modelling, featuring detailed schema annotations. It covers services like flight booking, calendar scheduling, banking, and media services, including unseen services at test time. We use the SGD test split for evaluation in the main results, containing 4,201 dialogues and 18 domains. We load these with ConvLab-3 Zhu et al. (2023).
General Ontology Data
To test whether TeQoDO’s concepts apply beyond TOD data, we apply it to the Wikipedia and ArXiv datasets by Lo et al. (2024). The Wikipedia test set contains 242,148 article titles and abstracts; its gold ontology graph has 4 hierarchy levels with 8,033 nodes and 14,673 edges. Topics include various types of “injuries” or “legislative bodies”.
The ArXiv test set has 27,630 title-abstract pairs, 2 hierarchy levels with 61 nodes and 61 edges. Main topics are ArXiv categories like “Mathematics” with subcategories such as “Commutative Algebra”, making this ontology more abstract and higher level than Wikipedia’s. These ontologies do not follow the domain-slot-value hierarchy resembling SQL’s table-column-value structure. TeQoDO is applied to the test sets of both datasets.
4.2 Evaluation
We run TeQoDO on a dataset, generating the ontology by executing all update queries per dialogue. For evaluation, tables map to domains, columns to slots, and values to ground truth values. System actions and user intents are matched by aligning table names with the “system actions” “and user intents” keys and comparing values directly to the ground truth.
We adapt the evaluation framework of Lo et al. (2024) to our task-oriented dialogue ontology structure, which includes domains, slots, values, system actions, and user intents. Specifically, we use literal F1 as a hard metric and fuzzy and continuous F1 as soft metrics, as these capture most relevant evaluation aspects Lo et al. (2024). The hard metric captures syntactic similarity, respectively, while the soft metrics capture some human intuition on semantic similarity. Since these metrics treat all ontology graph edges equally — overlooking infrequent higher-level relations — we extend them to account for hierarchical levels and their specific edge types.
We compute the macro average of each metric across the node classes: domains, slots, values, intents, and actions. Let $V_{i_{\text{{pred}}}}$ denote the predicted nodes and $V_{i_{\text{{true}}}}$ the ground truth for class $i$.
Literal Metric
In the literal score, only exact term matches count: true positives are $V_{i_{\textup{pred}}}\cap V_{i_{\textup{true}}}$, false positives are $V_{i_{\textup{pred}}}\setminus V_{i_{\textup{true}}}$, and false negatives are $V_{i_{\textup{true}}}\setminus V_{i_{\textup{pred}}}$.
Fuzzy Metric
In both the fuzzy and continuous setups, we map nodes above a similarity threshold to ground truth nodes. We use a sentence transformer and set the threshold at $t_{\textup{{sim}}}=0.436$ to consider two concepts equivalent as in (Lo et al., 2024). For the fuzzy setup, all predicted nodes above the threshold are mapped to ground truth. True positives are predicted nodes with cosine similarity above the threshold to at least one ground truth node, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\}$. False positives are predicted nodes with similarity below or equal to the threshold for all ground truth nodes, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\forall v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$. False negatives are ground truth nodes without any predicted node mapped above the threshold, i.e., $\{v_{t}\in V_{i_{\textup{true}}}\mid\forall v_{p}\in V_{i_{\textup{pred}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$.
Continuous Metric
For the continuous metric, only the predicted node with the highest similarity to each ground truth node above $t_{\textup{sim}}$ is a true positive, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\,|\,\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\land\operatorname{sim}(v_{p},v_{t})=\max_{v_{p}^{\prime}\in V_{i_{\textup{pred}}}}\operatorname{sim}(v_{p}^{\prime},v_{t})\}.$ This stricter metric penalises multiple predictions for the same ground truth node. We use this as the primary metric, as it accounts for surface-form variations without allowing significantly different structures — an approach that, through qualitative analysis, proved to reward the best ontologies in terms of downstream usability.
We match the hierarchy top-down: domains first, then slots, values, intents, and actions. Only slots of matched domains are considered; unmatched domains exclude their slots. This applies similarly to values, intents, and actions.
On Wikipedia and ArXiv, comparison results are from Lo et al. (2024) using their evaluation scripts. We report their literal, fuzzy, continuous, and graph F1 metrics. Graph F1 embeds predicted and ground truth graphs, comparing their structure through graph convolutions.
4.3 Models
DORE (Vukovic et al., 2024)
stands for dialogue ontology relation extraction, which uses Gemma-2B instruct (Team et al., 2024) to predict dialogue-level ontology relations between known terms in the prompt. DORE enhances transfer learning via constrained chain-of-thought decoding (Wang and Zhou, 2024), restricting decoding to known terms and relations. Final outputs are chosen based on confidence across sampled responses. Table˜1 presents the best-performing DORE models on both TOD datasets, including fine-tuned and transfer models, trained on the respective other dataset.
GenDSI (Finch et al., 2024)
is the generative dialogue state inference approach (GenDSI). It fine-tunes T5-3B (Raffel et al., 2020) on a large synthetic TOD dataset (Finch and Choi, 2024) to generate dialogue state updates from user-system turns. State updates, as slot-values, are clustered, with the most frequent slot name in each cluster chosen as representative. For evaluation, domains are extracted from slots using the first word of each slot name to form the ontology hierarchy. Predictions lacking domain names are discarded. As DORE and GenDSI do not predict user intents or system actions, evaluation is limited to domains, slots, and values, with averages calculated accordingly in Table˜1.
Approach Domains Slots Values Intents Actions Average Supervised
MultiWOZ
TeQoDO (ours) 60.04 57.25 70.65 56.21 82.11 65.25
DORE fine-tuned on MWOZ 15.53 20.39 84.01 - - 39.98
DORE fine-tuned on SGD 15.24 2.94 22.14 - - 13.44
GenDSI 29.79 25.64 33.94 - - 29.15
SGD
TeQoDO (ours) 72.19 43.70 48.57 76.48 67.28 61.64
DORE fine-tuned on MWOZ 12.84 16.35 60.04 - - 29.74
DORE fine-tuned on SGD 17.27 11.94 85.05 - - 38.05
GenDSI 27.07 29.08 41.28 - - 32.47
Table 1: Ontology construction comparison. DORE (Vukovic et al., 2024) and GenDSI (Finch et al., 2024) do not predict the intents and actions; hence, their average is only over domains, slots, and values.
OLLM (Lo et al., 2024)
is fine-tuned on Wikipedia and ArXiv ontologies and uses a custom masking loss and rule-based aggregation of article-level predictions. On these datasets, we also compare to Hearst patterns Roller et al. (2018), which use hand-crafted lexico-syntactic rules, e.g. “mammals such as humans”, to predict is-a relations. REBEL Huguet Cabot and Navigli (2021) treats relation extraction as a translation task and trains a language model to generate relations.
TeQoDO
We use GPT-4o-mini (“gpt-4o-mini-2024-07-18”) OpenAI (2024) for all experiments. The model performs well on HumanEval Du et al. (2024), which includes SQL coding tasks. Full pipeline prompts appear in Section˜A.1. SQLite [^3] https://www.sqlite.org is used via Python’s sqlite3.444https://docs.python.org/3/library/sqlite3.html Final prompts were manually crafted according to best practices [^5] https://help.openai.com/en/articles/6654000-best-practices-for-prompt-
engineering-with-the-openai-api
and refined with ChatGPT (see Section˜B.2 for prompt variant performance).
For more consistency, we sample values for each table column to expose the LLM to concrete instances in the column value example set-up.
To reduce variation in column values (e.g. “Alexander Bed and Breakfast” vs “Alexander B & B”) and improve naming consistency in the DB, we also experiment with query results for similar tables, columns, and values. We call this similarity matching. This decreases duplicate entries by increasing the chance of reusing existing values. The model tends to use the LIKE operator in SQLite, which matches substrings but fails in cases like the example above. Instead, we compute semantic similarity using the “all-MiniLM-L6-v2” sentence-transformers model Reimers and Gurevych (2019) and a similarity threshold of 0.4360.4360.436. This threshold, found by Lo et al. (2024), corresponds to the median similarity for synonyms in WordNet Miller (1994) using the same model. For each query, we return up to 5 similar concepts that exceed this threshold.
4.4 Task-oriented Ontology Construction
SOTA Comparison
Table˜1 shows that DORE overfits to domains and slots in the training set but outperforms TeQoDO on value-level and literal metrics. This stems from supervised fine-tuning, enabling DORE to learn exact value phrasing. GenDSI surpasses DORE on domains and slots but is outperformed by TeQoDO on all metrics. GenDSI’s lack of explicit domain prediction causes slot predictions to be dropped if domains are not the first word in slot names. These results show that TeQoDO predicts higher-level hierarchical concepts far better than the other methods.
Ablation
Table˜2 shows the ablation study results, while Section˜B.1 details results for domains, slots, values, intents, and actions.
In the table, direct update refers to generating updates without querying the database. Query update involves querying the database before applying updates. DST Step indicates the use of the DST step. Success indicates that user goal fulfilment is included in the update prompt (Section˜3.2). Sim. denotes the use of similarity matching. Ex. refers to column value examples (Section˜4.3).
The large discrepancy between literal and similarity-based metrics aligns with the findings of Lo et al. (2024), highlighting numerous surface form variations that literal evaluation fails to capture. For example, a domain named “hotel_bookings” is a false positive in literal evaluation despite being a reasonable prediction, since the ground truth domain is “hotel”. Comparing fuzzy and continuous metrics, precision on the latter is lower on both datasets, as only one of several predictions can match each ground truth concept.
Using the query and update pipeline improves all scores on both datasets, already surpassing DORE in the more expressive similarity-based metrics Lo et al. (2024). This is expected, as allowing the model to query existing tables leads to better alignment and more consistent update queries. Without querying first, the model tends to create new table names for every dialogue, reducing precision. Recall is generally higher in the direct update approach due to generating more tables, but this substantially lowers precision.
Approach Literal Fuzzy Continuous
F1 Precision Recall F1 Precision Recall F1 Precision Recall
MultiWOZ
Direct Update $2.8_{\pm 0.2}$ $2.5_{\pm 0.3}$ $16.1_{\pm 0.8}$ $37.7_{\pm 3.2}$ $32.2_{\pm 2.3}$ $88.1_{\pm 0.5}$ $19.2_{\pm 1.0}$ $19.9_{\pm 0.4}$ $58.5_{\pm 9.5}$
Query Update $13.1_{\pm 2.4}$ $14.7_{\pm 2.2}$ $13.6_{\pm 2.7}$ $62.3_{\pm 9.2}$ $60.7_{\pm 7.9}$ $77.7_{\pm 1.6}$ $50.1_{\pm 6.7}$ $45.0_{\pm 6.3}$ $73.4_{\pm 6.2}$
+ DST Step $11.5_{\pm 3.5}$ $14.5_{\pm 4.5}$ $12.2_{\pm 4.0}$ $70.4_{\pm 4.7}$ $70.3_{\pm 6.2}$ $75.5_{\pm 3.8}$ $58.8_{\pm 4.3}$ $54.5_{\pm 4.9}$ $74.8_{\pm 3.8}$
+ DST, Sim. $11.1_{\pm 3.1}$ $12.5_{\pm 4.4}$ $12.7_{\pm 3.7}$ $71.3_{\pm 3.1}$ $69.4_{\pm 4.8}$ $77.7_{\pm 2.9}$ $\mathbf{60.7}_{\pm 3.7}$ $55.0_{\pm 4.9}$ $77.1_{\pm 2.9}$
+ DST, Ex. $12.1_{\pm 1.8}$ $16.4_{\pm 2.5}$ $12.4_{\pm 2.3}$ $\mathbf{74.7}_{\pm 4.0}$ $74.3_{\pm 5.9}$ $79.8_{\pm 2.2}$ $\mathbf{65.2}_{\pm 5.5}$ $61.2_{\pm 6.5}$ $78.5_{\pm 2.4}$
+ DST, Ex., Success $12.3_{\pm 4.3}$ $16.8_{\pm 7.8}$ $12.9_{\pm 3.4}$ $73.3_{\pm 4.6}$ $75.3_{\pm 5.9}$ $76.2_{\pm 5.0}$ $\mathbf{65.2}_{\pm 4.7}$ $64.6_{\pm 7.0}$ $75.3_{\pm 4.7}$
SGD
Direct Update $2.5_{\pm 0.2}$ $2.1_{\pm 0.3}$ $17.4_{\pm 0.2}$ $42.6_{\pm 7.0}$ $34.9_{\pm 7.5}$ $91.2_{\pm 0.7}$ $18.6_{\pm 2.3}$ $18.6_{\pm 1.6}$ $66.7_{\pm 13.0}$
Query Update $7.0_{\pm 2.7}$ $13.0_{\pm 5.5}$ $6.5_{\pm 3.4}$ $53.8_{\pm 24.2}$ $76.9_{\pm 20.3}$ $54.5_{\pm 28.1}$ $44.6_{\pm 21.1}$ $64.0_{\pm 19.7}$ $49.2_{\pm 26.6}$
+ DST Step $10.0_{\pm 1.9}$ $15.7_{\pm 4.4}$ $8.9_{\pm 1.8}$ $75.0_{\pm 2.1}$ $89.1_{\pm 8.9}$ $69.1_{\pm 7.3}$ $61.4_{\pm 9.1}$ $72.2_{\pm 16.3}$ $61.9_{\pm 10.3}$
+ DST, Sim. $7.0_{\pm 2.6}$ $11.8_{\pm 5.7}$ $6.5_{\pm 1.9}$ $70.1_{\pm 5.1}$ $85.5_{\pm 9.2}$ $65.2_{\pm 10.7}$ $53.4_{\pm 7.4}$ $65.6_{\pm 14.4}$ $55.1_{\pm 12.4}$
+ DST, Ex. $9.0_{\pm 1.9}$ $14.7_{\pm 4.8}$ $8.2_{\pm 2.5}$ $73.5_{\pm 3.2}$ $90.3_{\pm 8.7}$ $66.1_{\pm 10.0}$ $60.8_{\pm 7.5}$ $74.3_{\pm 17.2}$ $59.5_{\pm 10.8}$
+ DST, Ex., Success $8.3_{\pm 2.2}$ $14.7_{\pm 5.1}$ $7.0_{\pm 2.8}$ $70.2_{\pm 5.7}$ $93.4_{\pm 6.2}$ $60.5_{\pm 10.5}$ $\mathbf{61.6}_{\pm 5.4}$ $80.7_{\pm 12.5}$ $58.0_{\pm 10.7}$
Table 2: TOD ontology construction ablation study for MultiWOZ and SGD test sets. We report the macro average scores and standard deviation for 5 seeds of different dialogue orders over the five hierarchy classes: domains, slots, values, system actions, and user intents. Bold F1 scores are significantly better than Query Update ($p<0.05$). On SGD Query Update approaches, we input batches of 10 dialogues for faster inference.
Approach # Tables SQL Error Ratio $\downarrow$
MultiWOZ (6 domains)
Direct Update 168 28.72%
Query and Update 14 37.31%
+ DST 12 30.45%
+ DST, Sim 16 23.19%
+ DST, Ex. 9 37.76%
+ DST, Ex. Success 9 4.71%
SGD (18 domains)
Direct Update 432 30.27%
Query and Update 92 32.50%
+ DST 69 32.81%
+ DST, Sim 88 17.71%
+ DST, Ex. 67 22.92%
+ DST, Ex. Success 37 5.22%
Table 3: Comparison of Approaches by Number of Tables and SQL Error Ratio in update queries.
Ontology Domains
hotel rest. attr. train taxi avg.
Ground truth Ontology 41.3 25.2 24.9 30.9 28.3 30.1
TeQoDO Ontology 26.7 23.9 44.8 29.1 33.2 31.5
Table 4: Zero-shot DST results for TripPy-R (Heck et al., 2022) with ground truth and TeQoDO induced ontology inference on MultiWOZ 2.1 in joint goal accuracy per domain.
On both datasets, incorporating dialogue theory improves performance and reduces variance. The DST step alone does not significantly outperform the query and update baseline. On MultiWOZ, similarity matching, column value examples, and success yield significantly better performance. On SGD, using success together with column value examples significantly improves the continuous metric. Note that the combination of similarity matching and success yields no improvements, and is therefore omitted for brevity. Dialogue theory narrows the gap between fuzzy and continuous metrics, indicating more consistent DB updates. The influence of dialogue order is notably diminished, as evidenced by the drop in variance when dialogue theory is applied.
Comparing the direct update baseline to the query and update pipeline, the number of tables is significantly reduced. As seen in Table˜3, the direct update baseline yields 168 tables for MultiWOZ, while the pipeline results in only 14, much closer to the ground truth of 6 domains. Pipeline variants using DST show similar table counts, which are further reduced by incorporating success, closest to the 18 ground truth domains. This reduction is even more pronounced on the SGD dataset. The direct update baseline produces many redundant tables per domain, e.g., over 15 train-related domains. In contrast, the pipeline typically results in one train_bookings table.
Incorporating dialogue success into the DB update prompt significantly reduces erroneous SQL query ratios, achieving single-digit error rates on both datasets. This suggests that applying dialogue theory improves the model’s handling of the database and generation of SQL update queries, as most errors arise from incorrect column names, indicating poor schema handling.
4.5 Downstream Application: DST
Approach Literal F1 Fuzzy F1 Continuous F1 Graph F1 Supervised
Wikipedia
TeQoDO 0.03 $\dagger$64.94 34.76 $\dagger$64.15
Hearst Patterns (Roller et al., 2018) 0.30 53.80 35.00 54.40
REBEL (Huguet Cabot and Navigli, 2021) $\dagger$0.40 62.40 $\dagger$35.60 7.20
OLLM (Lo et al., 2024) 9.30 91.50 50.00 64.40
ArXiv
TeQoDO + Clustering 0.00 $\dagger$33.95 $\dagger$34.15 89.89
Hearst Patterns (Roller et al., 2018) 0.00 0.00 15.10 55.30
REBEL (Huguet Cabot and Navigli, 2021) 0.00 6.00 28.10 54.60
OLLM (Lo et al., 2024) 4.00 57.00 35.70 $\dagger$63.30
Table 5: Ontology construction results for Wikipedia and ArXiv test sets from Lo et al. (2024). We use the metrics from their code base and the Hearst, REBEL and OLLM results from their work. In bold the best F1 score for each column is highlighted and with a $\dagger$ the second highest.
To demonstrate downstream usability, we apply the TeQoDO-induced ontology in a specialised dialogue state tracking model. We replicate the zero-shot leave-one-domain-out setup from Heck et al. (2022). Using their TripPy-R model, we train with one domain held out and infer on that domain using the TeQoDO-induced ontology instead of ground truth. Only domain and slot predictions are used in this setup.
See Table˜4 for results comparing inference using the ground truth and TeQoDO-induced ontology. TripPy-R performs similarly with the induced ontology across all domains except “hotel”. In this case, the slots hotel-book people and hotel-book stay are missing due to unmapped slot predictions. The “restaurant” and “train” domains also lack some slots, leading to lower performance. In the “taxi” and “attraction” domains, all slots are mapped, and performance surpasses that with ground truth slots. This may result from more informative slot names in the induced ontology, e.g. taxi_bookings-pickup_location instead of ground truth taxi-departure. This shows that TeQoDO can induce a useful ontology for downstream DST.
4.6 General Ontology Construction
When applying TeQoDO to general ontologies, we truncate the prompt from the left to fit the model’s context window, needed only for Wikipedia. Wikipedia’s table count reaches thousands, indicating that pre-filtering tables in TeQoDO’s first step may be necessary—this is left for future work.
On ArXiv, we evaluate only the first two hierarchy levels, as the ground truth ontology contains just two. Due to the dataset’s size and the small target ontology, directly applying TeQoDO results in too many predicted tables. To address this, we cluster table names and select representative tables based on their proximity to each centroid. This clustering approach mirrors OLLM’s frequency-based aggregation but avoids hand-crafted rules. We use $k$-means MacQueen et al. (1967), choosing $k$ via the silhouette score Rousseeuw (1987), with $k$ ranging from 5 to 20. To speed up inference, we prompt with batches of 5 articles for ArXiv and 200 for Wikipedia.
Results
See Table˜5 for the final TeQoDO results compared to models from Lo et al. (2024) (see Section˜A.4 for prediction and ground truth examples). On Wikipedia, TeQoDO does not outperform OLLM on most metrics, likely due to OLLM’s extensive supervised training. However, TeQoDO surpasses REBEL in fuzzy and graph F1, where it performs comparably to OLLM. On the continuous metric, TeQoDO performs on par with Hearst patterns and REBEL.
By clustering and merging tables, we reduce predicted table numbers and achieve competitive results on ArXiv. TeQoDO attains the highest graph F1 and second-best continuous F1 on ArXiv, outperforming fine-tuned and few-shot models and matching supervised OLLM. This shows TeQoDO’s induced ontologies closely match the ground truth structure. In fuzzy and continuous F1, TeQoDO surpasses Hearst patterns and REBEL. Lower literal F1 scores reflect the absence of supervision in TeQoDO. Results suggest text-to-SQL can adapt to ontologies beyond TOD, though ArXiv requires adjustments. The differing hierarchy depth poses challenges, as SQL suits the three-level TOD structure best. We argue that ArXiv’s high-level labels, with only 61 nodes for thousands of abstracts, underrepresent its content richness.
Query Update Query Update + DST
INSERT INTO hotel_bookings (hotel_name, location, price_category, star_rating) VALUES (‘Alexander Bed and Breakfast’, ‘Centre’, ‘cheap’, 4); UPDATE hotel_details SET address = ‘56 saint barnabas road’, phone_number = ‘01223525725’ WHERE name = ‘Alexander Bed and Breakfast’ AND address IS NULL AND phone_number IS NULL;
Query Update Query Update + DST Similarity Matching
SELECT name, location, free_wifi FROM guesthouses WHERE name = ‘Allenbell’;
$\rightarrow$ No Result
SELECT name, location, free_wifi FROM guesthouses WHERE name = ‘The Allenbell’;
$\rightarrow$ Result: [(‘The Allenbell’, ‘east’, 0)]
Query Update Query Update + DST Column Value Examples
INSERT INTO intents (intent) VALUES (’find_swimming_in_east’); INSERT INTO intents (intent_name) VALUES (’find_pool’);
Query Update Query Update + DST Value Examples and Success
INSERT INTO actions (action) VALUES (’provide_information’), (’recommend’); INSERT INTO system_actions (action) VALUES (’ask_clarification’);
Table 6: Qualitative comparison between the query and update baseline and proposed improvements. Note that the hotel tables were named differently in the different runs of the ablations.
The main challenges for TeQoDO on general ontologies are the dataset size and differing hierarchy levels compared to TOD ontologies. However, the graph F1 results indicate that the structure can be effectively induced on much larger ontologies, showing the generalisability of our approach.
4.7 TOD Ontology Construction Qualitative Analysis
When prompting the model to generate an SQL DB, the induced hierarchy aligns more closely with the desired structure than when explicitly prompting for domains, slots, and values. The latter leads to more misclassifications, such as predicting slots as domains.
See Table˜6 for a qualitative comparison of update queries on MultiWOZ test dialogues, detailed in Section˜A.3. In the similarity matching example, only “Allenbell” appears in the dialogue, while “The Allenbell” is stored from a previous dialogue. Without similarity matching, the model queries “Allenbell”, yielding no result due to the missing “The”. Similarity matching retrieves “The Allenbell” from the database. The model chooses which results to include, as less related concepts may also be retrieved.
In the second example, we illustrate the impact of adding the DST step, which clarifies the distinction between new dialogue information and existing DB content. Without DST, the model inserts a new entry for “Alexander Bed and Breakfast”. With DST, it updates the existing DB entry by adding only the missing information.
Column value examples help the model generate more general intents and actions, e.g., find_pool” instead of “find_swimming_in_east”. Mentioning dialogue success encourages system actions like “ask_clarification” that support user goals. See Section˜A.2 for predicted ontology excerpts from the “restaurant” (MultiWOZ) and “flights” (SGD) tables.
4.8 Discussion and Future Work
Our results show that SQL’s structured format improves ontology quality when LLMs interact with a database. The SQL-enhanced model aligns values more accurately using semantic matching or column value examples. Dialogue theory distinguishes existing database information from new input, making the database more user-focused. TeQoDO significantly outperforms the supervised fine-tuned DORE and GenDSI. It also reduces sensitivity to dialogue order. Future work will enable database restructuring through a global view to minimise order effects, since the model could unify table names instead of clustering, as done on ArXiv. Smaller models could improve scalability.
Regarding training data contamination, the results of the direct update approach—with its large variance in table naming—suggest that the model has not memorised the TOD datasets, as it cannot recall the exact table names for each dialogue
For larger or structurally different ontologies than TOD, our results show TeQoDO can generalise, though SQL struggles to express multiple hierarchy levels. We aim to scale TeQoDO to diverse ontology structures by adapting the text-to-SQL component; however, its sequential nature limits parallelisation. The large batch size for concatenated Wikipedia articles may affect performance, though evaluating this is computationally expensive. Finally, our approach is sensitive to prompt phrasing, a common issue with large language models Razavi et al. (2025).
Our evaluation captures all necessary information and is more fine-grained than existing methods by considering the hierarchy levels of TOD ontologies: domains, slots, values, system actions, and user intents. Using an off-the-shelf similarity model with a fixed threshold may introduce bias, though qualitative analysis confirms alignment with ontology quality. In future work, we aim to incorporate human judgement into evaluation to reduce reliance on fixed thresholds.
Dennett (1987) distinguishes competence from comprehension, arguing that the latter is not required for the former. Inspired by this, we view LLMs’ ability to perform tasks like SQL generation as a form of competence to extract task knowledge in the form of an ontology.
We see this work as an important step in distilling human-readable knowledge from otherwise black-box LLMs. Extracting ontologies automatically with LLMs and analysing their structure may enhance their interpretability in downstream tasks. For instance, we might hope to detect hallucinations based on the constructed ontology.
5 Conclusion
We introduce TeQoDO, a method that uses large language models to build ontologies from task-oriented data, exploiting the inherent SQL programming abilities of LLMs and incorporating dialogue modelling theory. We evaluate TeQoDO on two widely used TOD datasets and find that each proposed step improves performance, surpassing current SOTA. We also show that TeQoDO generalises to large ontologies with different structures. Our results motivate further exploration of this approach for explainability.
Appendix A Supplementary Information
A.1 TeQoDO Prompt
  1. 1.
    You’re working with a dialogue system that stores structured data from conversations in an SQLite3 database. The database includes tables covering user intents, system actions, and information about various entities. You will be provided with two inputs: the current set of database tables and their names (but not their schemas). New dialogue(s) – this contains user and system turns with references to specific intents, actions, or queried information. Your task:
    Identify which tables from {db_result_input} are relevant to the new dialogue(s) based on: User intents expressed in the dialogue(s). System actions performed in response. Specific information or entities being queried or discussed. For each relevant table, generate the following SQLite command to inspect its schema: PRAGMA table_info(<table_name>); This will allow you to understand the structure (columns and data types) of the tables you will be working with. Do not create or modify tables yet — only inspect existing ones using PRAGMA.
  2. 2.
    You’ve already examined the database schema using PRAGMA table_info(…) queries. The schema details are provided, which contain the structure (columns and types) of the current tables in the SQLite3 database. Now, based on the same dialogue and the table definitions, your task is to: Generate SQL SELECT queries to retrieve: User intents expressed in the dialogue — in general form (e.g., find_flight, book_hotel) without including specific parameters (e.g., cities, dates). Use the table column schema to determine which table holds this information and write a query to retrieve matching intents: {db_result_input}
    System actions carried out in the dialogue — again in a generalized form (e.g., recommend, confirm, inform). Use the appropriate table from above and generate a query to retrieve those action types. Information explicitly requested by the user — such as facts about entities (e.g., list of Italian restaurants, hotel prices, flight times), but only if that data is already present in the database. Generate SELECT queries from the relevant tables, based on what was asked in the dialogue. Do not: Create or alter tables (no CREATE, INSERT, or UPDATE); Use timestamp fields or session-specific filters; Use specific slot values from the dialogue (e.g., exact restaurant names or locations) in the intent or action queries — keep them general.
  3. 3.
    You’ve already run a set of SELECT queries based on the previous dialogue, and the results of those queries are provided in the following. These results represent all the information currently stored in the database that matches the dialogue. {db_result_input} Your task now is to perform Dialogue State Tracking (DST) by extracting structured information from the dialogue — but only if that information is already present in the database, as confirmed by the query results.
    Specifically: Use the dialogue to identify user intents, system actions, and information about entities (e.g., preferences, attributes, categories). For each matching element, only include it in the tracked state if it appears in the DB results above. Represent the extracted state using a table → column → value structure, reflecting exactly how the information maps to the current database. Do not: Track or infer values that do not exist in the current DB results; Use placeholder values or hypothetical interpretations; Modify, insert, or extend the database structure in any way. Your output should reflect only what is both: Mentioned in the dialogue, and Already stored in the database results above.
  4. 4.
    You have already reviewed the current database contents using SELECT queries. The structure and current entries of the database are given in: {db_result_input}. You’ve also performed Dialogue State Tracking (DST), which revealed the information from the dialogue that is already present in the DB. Now, based on the dialogue and what is missing from the DST results (i.e., what’s not yet in the DB), generate SQL queries (SQLite3 syntax) to bring the database up to date. This will ensure that the dialogue can be successfully handled using only the data in the database.
    Your SQL queries should: Insert missing user intents into the database using general labels (e.g., book_train, find_hotel), as observed in the dialogue and not yet stored according to DB results above. Insert missing system actions in generalized form (e.g., inform, offer_options, confirm_request) that were present in the dialogue but missing from the DB. Insert or update entity information: Use INSERT statements if an entity mentioned in the dialogue does not yet exist in the DB. Use UPDATE statements if an entity exists but is missing column values (e.g., NULL) that are provided in the dialogue. Modify the schema if needed: Use ALTER TABLE if the dialogue introduces a new attribute not present in any table based on the current DB results. Use CREATE TABLE if a new entity type is mentioned in the dialogue that has no table yet. Do not: Generate Python code — write only raw SQL queries; Use timestamps or session data; Update values that are already correctly populated; Add user intents or actions with specific slot values from the dialogue — keep them generalized. Once the updates are applied, the database should fully support executing and resolving the dialogue based on its contents, so that the user’s goal expressed in the dialogue can be successfully fulfilled using only information stored in the database. Make sure to have tables for different types of entities, e.g. a restaurants table, etc. and not one table for all entities.
Figure 4: Prompts for TeQoDO steps. db_result_input is the result of the DB queries from the prior step.
Baseline: ${}\mathcal{C}_{i}=L(d_{i}),\quad\forall d_{i}\in\mathcal{D},\quad\mathcal{O}_{i}=\mathcal{C}_{0}\cup\dots\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Iterative Baseline: ${}\mathcal{C}_{i}=L(d_{i},\mathcal{O}_{i-1}),\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Tracking: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i}),\quad\mathbf{b}_{i}=$ $\mathcal{O}_{i-1}(d_{i})=L(d_{i},\mathcal{O}_{i-1},L(d_{i},\mathcal{O}_{i-1}))$ , ${}\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Success: $\mathcal{C}_{i}=L(d_{i},\mathbf{b}_{i})$,  ${}\mathbf{b}_{i}=\mathcal{O}_{i-1}(d_{i},\text{is\_successful}(d_{i})),\quad\mathcal{O}_{i}=\mathcal{O}_{i-1}\cup\mathcal{C}_{i},\quad\mathcal{O}_{0}=\{\varnothing\}$
Table 7: Update formulas for different SQL-based ontology construction prompts. $\mathcal{C}_{i}$ are the update messages for a dialogue in structured language $L$, where we use SQL. The dialogues in the dataset are $\mathcal{D}=\{d_{0},\dots,d_{n}\}$. The ontology after dialogue $d_{i}$ is $\mathcal{O}_{i}$ and $\mathbf{b}_{i}$ is the belief state. The final ontology contains $m$ concepts: $\mathcal{O}=\{c_{1},\dots,c_{m}\}$. $\mathbf{b}_{i}=L(d_{i},\mathcal{O}_{i-1})$ is the domain-slot information extracted from the DB for the current dialogue.
A.2 Predicted Ontology Excerpts
restaurants’: {’food_type’: {’african’,
asian’,
asian oriental’,
brazilian’,
british’,
chinese’,
european’,
french’,
gallery’,
gastropub’,
general’,
guesthouse’,
indian’,
international’,
italian’,
japanese’,
korean’,
mediterranean’,
modern european’,
portuguese’,
seafood’,
spanish’,
thai’,
name’: {’acorn guest house’,
anatolia’,
ask restaurant’,
bangkok city’,
bloomsbury restaurant’,
cafe jello gallery’,
caffe uno’,
cambridge chop house’,
cambridge lodge’,
charlie chan’,
chiquito’,
city stop’,
city stop restaurant’,
phone_number’: {’01223241387’,
’01223308681’,
’01223312112’,
’01223351707’,
’01223354755’,
’01223358899’,
’01223362433’,
’01223365599’,
’01223368786’,
price_range’: {’cheap’,
expensive’,
moderate’,
moderately priced’,
pricey’,
varied’}}
Figure 5: JSON representation of restaurant table predicted via TeQoDO on MultiWOZ test set.
flights’: {’airline’: {’air france’,
alaska airlines’,
american airlines’,
delta airlines’,
generic airline’,
n/a’,
new airline’,
placeholder airline’,
southwest airlines’,
united airlines’},
arrival_city’: {’arrival city’,
atlanta’,
atlanta, ga’,
berlin’,
chi-town’,
chicago’,
london’,
los angeles’,
nairobi’,
new york’,
new york city’,
none’,
nyc’,
paris’,
philadelphia’,
philly’,
phoenix’,
portland’,
portland, or’,
rio de janeiro’,
san diego’,
san francisco’,
seattle’,
sf’,
sydney’,
toronto’,
vancouver’,
washington’},
arrival_time’: {’01:50 pm’,
n/a’,
none’,
tbd’},
class’: {’n/a’, economy’, premium economy’},
departure_city’: {’atl’,
Figure 6: JSON representation of restaurant table predicted via TeQoDO on SGD test set.
A.3 Example Dialogues
See Figure˜7 for the example input dialogues from MultiWOZ test set for the generated queries in Table˜6.
Dialogue for Similarity Matching Example (multiwoz-test-300)
User: I am looking for a cheap place to stay.
System: We have ten cheap options, any preferences.
User: We will need free parking. Do any of the options have 4 stars?
System: Allenbell matches your requirements, would that work for you?
User: I need the hotel type, address and to know if they have wifi.
System: Allenbell is a guesthouse, located at 517a Coldham Lane. Yes, they do have wifi.
User: I’m also looking for a particular attraction. It’s called Trinity College.
System: Trinity College’s phone number is 01223338400. They are on Trinity Street with postcode CB21TQ.
User: Thank you so much for the phone number. Goodbye.
System: You’re welcome. Have a nice day.
Dialogue for DST Step Example (multiwoz-test-100)
User: Hi, I’m looking for some places to go in Cambridge.
System: Sure. Do you know what you would like to do?
User: A park or swimming on the East side.
System: There is Cherry Hinton Park and Abbey Pool and Astroturf Pitch. Would you like more information?
User: What is the phone number for Abbey Pool and Astroturf Pitch?
System: The phone number is 01223902088. Can I help with anything else?
User: Could you also tell me Abbey Pool’s address?
System: The address is Pool Way, Whitehill Road, off Newmarket Road. Can I help you any further?
User: I would also like help finding a guesthouse.
System: What type of guesthouse are you looking for?
User: I’d like to find a cheap 4-star place in the cheap price range, please.
System: Which area would you like to stay in?
User: The area does not matter. Just whatever you think is best in that criteria.
System: In that case, I would recommend the Alexander Bed and Breakfast in the centre.
User: Thank you! Do they have free WiFi?
System: Yes, they do. Would you like me to book a reservation for you?
User: No, thanks. I just need their address and phone number.
System: Sure thing. Their phone number is 01223525725. The address is 56 Saint Barnabas Road.
Anything else I can do for you?
User: No, that will be all. Thank you!
System: You’re very welcome! Take care!
Figure 7: Example dialogues for Table˜6 queries.
A.4 General Ontology Prediction Examples
Dataset Predicted Example SQL Queries Ground truth Edge Excerpt
ArXiv
CREATE TABLE IF NOT EXISTS Mathematics (
    id INTEGER PRIMARY KEY,
    category_name TEXT,
    parent_id INTEGER
INSERT INTO Mathematics (category_name, parent_id)
VALUES (’Statistical Analysis Techniques’, 1);
(Quantitative Finance, Risk Management), (Quantitative Finance, Statistical Finance), (Quantitative Finance, Trading and Market Microstructure), (Statistics, Applications), (Statistics, Computation), (Mathematics, Commutative Algebra), (Mathematics, Algebraic Geometry), (Mathematics, Statistics Theory)
Wikipedia
CREATE TABLE Theatre (
    id INTEGER PRIMARY KEY,
    category_name TEXT,
    parent_id INTEGER
CREATE TABLE Performance (
    id INTEGER PRIMARY KEY,
    category_name TEXT,
    parent_id INTEGER
INSERT INTO Theatre (category_name, parent_id)
VALUES (’Dialect Coaching’, 1);
INSERT INTO Theatre (category_name, parent_id)
VALUES (’Diction Coaching’, 1);
INSERT INTO Theatre (category_name, parent_id)
VALUES (’Dramaturgy’, 1);
INSERT INTO Theatre (category_name, parent_id)
VALUES (’Costume Design’, 1);
INSERT INTO Performance (category_name, parent_id)
VALUES (’Live Performance’, 1);
INSERT INTO Performance (category_name, parent_id)
VALUES (’Theatrical Performance’, 1);
(Injuries, Wounded and disabled military veterans topics), (Injuries, Healing), (Local government, Seats of local government), (Local government, Unincorporated areas), (Local government, Water management authorities), (Scientific problems, Unsolved problems in astronomy), (Youth health, Sex education), (History of organizations, History of schools), (Design history, Architectural history), (Theatrical occupations, Acting), (Theatrical occupations, Dance), (Theatrical occupations, Dance occupations), (Theatrical occupations, Diction coaches), (Theatrical occupations, Drama teachers)
Table 8: Examples comparing predicted SQL queries to ground truth edges, categorised by dataset.
Appendix B Complementary Results
B.1 TOD Ontology Results per Nodeclass
Approach EvalType Literal Fuzzy Continous
F1 Precision Recall F1 Precision Recall F1 Precision Recall
Direct Update Baseline Micro 1.15 0.94 1.49 63.57 56.75 72.24 32.29 23.97 49.46
Macro 2.96 3.27 16.34 37.16 31.50 88.92 19.50 19.75 62.96
Domains 6.29 3.31 62.50 21.21 12.07 87.50 8.70 4.58 87.50
Slots 6.17 3.73 17.86 42.74 28.09 89.29 18.40 11.11 53.57
Values 2.35 9.29 1.35 80.17 93.48 70.17 59.75 77.43 48.64
User intents 0.00 0.00 0.00 11.87 6.32 98.59 3.16 1.62 69.01
System actions 0.00 0.00 0.00 29.81 17.55 99.02 7.50 4.02 56.10
Iterative Query and Update Baseline Micro 13.76 26.03 9.35 78.03 85.58 71.71 65.84 69.44 62.60
Macro 15.55 18.11 16.26 69.01 65.06 79.25 47.33 42.83 61.42
Domains 47.06 44.44 50.00 62.50 62.50 62.50 55.56 50.00 62.50
Slots 15.38 12.00 21.43 60.61 52.63 71.43 49.35 38.78 67.86
Values 15.28 34.11 9.85 78.37 89.08 69.97 69.37 75.19 64.39
User intents 0.00 0.00 0.00 53.97 37.57 95.77 43.45 28.10 95.77
System actions 0.00 0.00 0.00 89.59 83.54 96.59 18.94 22.08 16.59
+ Similarity Matching Micro 17.71 34.88 11.87 75.22 89.90 64.66 70.97 79.01 64.42
Macro 20.56 27.72 19.29 71.86 71.01 79.24 61.02 56.84 79.19
Domains 62.50 62.50 62.50 80.00 85.71 75.00 70.59 66.67 75.00
Slots 19.67 18.18 21.43 63.33 59.38 67.86 59.37 52.78 67.86
Values 20.62 57.89 12.54 75.32 95.17 62.32 73.81 91.03 62.06
User intents 0.00 0.00 0.00 55.46 39.52 92.96 30.41 18.18 92.96
System actions 0.00 0.00 0.00 85.17 75.28 98.05 70.90 55.52 98.05
+ Column Value Examples Micro 14.04 31.19 9.06 77.78 92.82 66.94 66.88 68.80 65.07
Macro 16.79 22.21 16.19 78.67 79.10 81.42 65.03 58.04 81.02
Domains 50.00 50.00 50.00 93.33 100.00 87.50 82.35 77.78 87.50
Slots 18.18 15.79 21.43 60.00 56.25 64.29 54.55 47.37 64.29
Values 15.75 45.26 9.54 77.18 95.39 64.81 67.17 72.22 62.79
User intents 0.00 0.00 0.00 71.74 58.41 92.96 43.42 28.33 92.96
System actions 0.00 0.00 0.00 91.12 85.47 97.56 77.67 64.52 97.56
+ DST Step Micro 11.16 21.80 7.50 73.57 92.61 61.02 69.47 82.16 60.18
Macro 9.58 13.10 8.74 72.38 74.91 74.44 60.96 59.26 73.55
Domains 23.53 22.22 25.00 66.67 71.43 62.50 58.82 55.56 62.50
Slots 11.76 13.04 10.71 65.38 70.83 60.71 59.26 61.54 57.14
Values 12.63 30.26 7.98 72.67 95.92 58.49 70.58 91.11 57.61
User intents 0.00 0.00 0.00 66.67 51.97 92.96 39.64 25.19 92.96
System actions 0.00 0.00 0.00 90.50 84.39 97.56 76.48 62.89 97.56
+ DST and Similarity Matching Micro 7.57 15.97 4.96 74.10 83.70 66.48 68.02 73.55 63.27
Macro 9.54 11.11 10.68 68.07 64.54 75.53 58.05 51.88 74.13
Domains 28.57 23.08 37.50 55.56 50.00 62.50 45.45 35.71 62.50
Slots 10.71 10.71 10.71 58.62 56.67 60.71 54.24 51.61 57.14
Values 8.40 21.75 5.21 73.64 85.93 64.42 68.52 78.20 60.97
User intents 0.00 0.00 0.00 64.08 48.89 92.96 44.30 29.07 92.96
System actions 0.00 0.00 0.00 88.44 81.22 97.07 77.73 64.82 97.07
+ DST and Column Value Examples Micro 3.84 12.13 2.28 82.18 95.55 72.09 78.19 88.54 70.01
Macro 13.15 17.84 11.52 81.86 85.80 81.46 72.15 71.00 81.01
Domains 42.86 50.00 37.50 85.71 100.00 75.00 80.00 85.71 75.00
Slots 18.87 20.00 17.86 78.43 86.96 71.43 72.73 74.07 71.43
Values 4.04 19.21 2.25 81.91 98.01 70.36 79.31 94.94 68.10
User intents 0.00 0.00 0.00 72.93 60.00 92.96 49.81 34.02 92.96
System actions 0.00 0.00 0.00 90.29 84.03 97.56 78.90 66.23 97.56
+ DST and Column Value Examples and Dialogue Success Micro 9.83 32.88 5.78 76.83 92.95 65.48 73.73 87.36 63.77
Macro 18.09 28.29 14.77 78.56 83.31 77.01 71.58 73.10 75.93
Domains 57.14 66.67 50.00 80.00 85.71 75.00 75.00 75.00 75.00
Slots 22.73 31.25 17.86 66.67 80.00 57.14 62.50 75.00 53.57
Values 10.56 43.53 6.01 75.78 94.26 63.36 73.30 90.61 61.54
User intents 0.00 0.00 0.00 75.86 64.08 92.96 57.89 42.04 92.96
System actions 0.00 0.00 0.00 94.51 92.52 96.59 89.19 82.85 96.59
Table 9: All classes results on MultiWOZ test set.
Approach EvalType Literal Fuzzy Continuous
F1 Precision Recall F1 Precision Recall F1 Precision Recall
Direct Update Baseline Micro 0.96 0.67 1.69 63.05 55.08 73.73 23.12 15.50 45.47
Macro 2.75 2.63 17.51 41.24 35.67 90.76 19.15 18.91 71.64
Domains 7.09 3.71 78.95 25.00 14.29 100.00 8.98 4.70 100.00
Slots 4.22 3.01 7.10 59.95 47.06 82.58 24.91 17.39 43.87
Values 2.45 6.43 1.51 81.65 94.98 71.60 52.11 67.45 42.46
User intents 0.00 0.00 0.00 17.73 9.73 99.60 4.77 2.45 89.33
Iterative Query and Update Baseline Micro 1.76 1.17 3.61 39.79 26.11 83.57 31.56 20.10 73.39
Macro 5.67 4.10 12.30 42.60 36.69 88.99 29.48 22.36 84.23
Domains 20.20 12.50 52.63 61.02 45.00 94.74 35.64 21.95 94.74
Slots 3.91 3.15 5.16 58.15 50.23 69.03 39.55 30.53 56.13
Values 4.22 4.87 3.72 82.55 82.40 82.70 61.54 53.83 71.81
User intents 0.00 0.00 0.00 5.48 2.82 98.81 5.15 2.64 98.81
System actions 0.00 0.00 0.00 5.83 3.00 99.66 5.53 2.84 99.66
+ Similarity Matching Micro 4.43 4.55 4.31 79.02 78.97 79.08 58.46 49.34 71.70
Macro 5.64 4.82 8.13 73.00 64.53 87.77 48.00 37.41 84.26
Domains 18.18 12.77 31.58 75.00 62.07 94.74 52.94 36.73 94.74
Slots 4.68 4.86 4.52 65.82 64.60 67.10 53.61 50.28 57.42
Values 5.34 6.46 4.55 80.03 82.41 77.79 62.82 57.04 69.91
User intents 0.00 0.00 0.00 66.05 49.51 99.21 32.26 19.26 99.21
System actions 0.00 0.00 0.00 78.11 64.09 100.00 38.38 23.75 100.00
+ Column Value Examples Micro 2.66 3.28 2.23 81.17 82.63 79.76 61.65 56.37 68.03
Macro 5.53 4.42 9.65 76.90 70.50 87.01 50.94 40.34 82.29
Domains 20.78 13.79 42.11 66.67 51.43 94.74 45.57 30.00 94.74
Slots 3.92 3.97 3.87 65.59 65.38 65.81 50.90 47.49 54.84
Values 2.97 4.31 2.26 81.17 83.82 78.69 64.07 62.20 66.06
User intents 0.00 0.00 0.00 84.92 75.62 96.84 45.58 29.81 96.84
System actions 0.00 0.00 0.00 86.13 76.23 98.99 48.60 32.21 98.99
+ DST Step Micro 4.50 5.93 3.63 80.22 85.98 75.17 59.53 56.35 63.10
Macro 6.69 6.05 11.38 74.47 69.61 84.67 47.55 38.88 79.12
Domains 21.69 14.06 47.37 72.00 58.06 94.74 41.86 26.87 94.74
Slots 6.57 7.56 5.81 63.31 71.54 56.77 43.62 45.45 41.94
Values 5.21 8.62 3.73 80.78 89.26 73.77 62.32 63.86 60.84
User intents 0.00 0.00 0.00 75.23 60.88 98.42 43.92 28.26 98.42
System actions 0.00 0.00 0.00 81.04 68.28 99.66 46.05 29.94 99.66
+ DST and Similarity Matching Micro 2.70 3.42 2.23 80.05 83.84 76.59 58.63 54.87 62.94
Macro 3.67 3.19 7.29 73.04 66.01 86.31 44.68 36.11 79.40
Domains 12.37 7.69 31.58 63.16 47.37 94.74 36.73 22.78 94.74
Slots 2.76 2.96 2.58 64.92 66.00 63.87 43.17 42.50 43.87
Values 3.22 5.27 2.31 80.60 86.86 75.19 62.38 64.22 60.64
User intents 0.00 0.00 0.00 74.55 60.00 98.42 40.06 25.15 98.42
System actions 0.00 0.00 0.00 81.99 69.81 99.33 41.05 25.87 99.33
+ DST and Column Value Examples Micro 3.36 4.53 2.67 81.52 85.40 77.98 62.95 59.08 67.37
Macro 7.16 6.23 12.35 76.59 71.02 86.53 50.06 41.07 81.04
Domains 24.69 16.13 52.63 69.23 54.55 94.74 42.86 27.69 94.74
Slots 7.38 8.62 6.45 67.81 72.26 63.87 49.33 51.03 47.74
Values 3.75 6.39 2.66 81.78 87.55 76.73 65.96 66.51 65.43
User intents 0.00 0.00 0.00 78.04 65.00 97.63 44.91 29.16 97.63
System actions 0.00 0.00 0.00 86.09 75.77 99.66 47.22 30.94 99.66
+ DST and Column Value Examples and Dialogue Success Micro 3.44 6.58 2.33 75.45 86.81 66.71 64.83 68.25 61.73
Macro 10.26 10.60 12.27 79.12 79.67 81.37 59.25 53.21 78.00
Domains 39.22 31.25 52.63 87.18 85.00 89.47 64.15 50.00 89.47
Slots 8.40 12.05 6.45 66.42 77.59 58.06 52.94 61.54 46.45
Values 3.70 9.70 2.29 74.78 88.78 64.60 66.21 74.82 59.38
User intents 0.00 0.00 0.00 78.39 66.21 96.05 52.77 36.38 96.05
System actions 0.00 0.00 0.00 88.82 80.77 98.66 60.18 43.30 98.66
Table 10: All classes results on SGD testset.
B.2 Prompt as Hyperparameter: Search Plots
See Figures˜8 and 9 for differences in performance for ChatGPT explanation-based prompts and success position in the prompts.
Figure 8: Continuous F1 Performance of different Prompts on MultiWOZ test-set based on LLM explanation.
Figure 9: Continuous F1 Performance of Prompts on MultiWOZ test-set with success mentioned in different steps.